OpenStudy (anonymous):

solve the equation: 8^z+1=32sqrt2

OpenStudy (anonymous):

I have to break this down into steps for the teacher

OpenStudy (nikvist):

\[8^z+1=32\sqrt{2}\quad\mbox{or}\quad 8^{z+1}=32\sqrt{2}\quad\mbox{???}\]

jhonyy9 (jhonyy9):

first write this in function of exponents of 2 like 8=2 on exponent 3 and 32=2 on exponent 5 sqrt2 will be 2 on exponent 1/2 - but you need to write again there is 8 on exponent z or z+1 ??? - so and after this you need using logaruthm on both sides and after you will can equlized those exponents from both sides getting hence .in this way one equation what you need to solve for z ... - so that's ALL